Circular No. 7529

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2000es, 2000et, 2000eu Further to IAUC 7524, L.-G. Strolger reports the discovery of three ad-

ditional supernovae by the Nearby Galaxies Supernova Search team. Direct confirmation of each object was made on Nov. 22 by B. Anthony-Twarog (CTIO 0.9-m telescope) and by Strolger (Kitt Peak 0.9-m telescope). Spectra (range 400–900 nm) were obtained by R. C. Smith and Strolger using the Mayall 4-m telescope on Nov. 22. From the spectra, N. B. Suntzeff and M. M. Phillips identify SN 2000es as a type-II event at z = 0.074. The spectra of SN 2000et indicate that it is a type-Ia event near maximum light, and SN 2000eu is a type-Ia event (possibly pre-maximum).

SN 2000 UT α_{2000} δ_{2000} R Offset
2000es Nov. 16 4^{h}06^{m}17.^{s}87 − 3^{o}23′18′′.2 19.7 4′′.6 W, 8′′ N
2000et Nov. 17 8 21 15.21 + 1 09 25.3 19.4 0′′.8 W, 5′′.2 N
2000eu Nov. 16 9 00 13.00 − 5 35 04.2 18.7 1′′.6 E, 1′′.2 S

SUPERNOVA 2000ev IN UGC 3500 M. Villi, Piacenza, Italy, reports the discovery of a supernova (mag

16.5) on an unfiltered CCD image taken on Nov. 27.10 UT by Federico Manzini (Soffrago, Italy) with a 0.40-m telescope. SN 2000ev is located 4′′.5 west and 19′′ north of the nucleus of UGC 3500 (α = 6^{h}46^{m}.9, δ = +84^{o}10′, equinox 2000.0), and it also appears on an image taken on Nov. 27.2 by R. Crippa (Tradate, Italy); nothing is visible at this position on images taken by Manzini on Sept. 1 or on the Digital Sky Survey.

C. Pernechele, S. Benetti, E. Cappellaro, S. Desidera, E. Giro, and L. Traverso, Astronomical Observatory of Padova, write: “Inspection of a preliminarily reduced spectrum (range 350–845 nm; resolution 0.2 nm) taken on Nov. 27 with the Asiago Ekar 1.82-m telescope (+ AFOSC) shows a very blue continuum with superimposed narrow Balmer emissions. The line profiles show two components, one unresolved and a broader one with FWHM of ∼ 2500 km/s. A faint He I 587.6-nm line is also present. This indicates that SN 2000ev is a type-IIn supernova. A polarization observation made with an R filter (50 min total exposure; standard polarized star 9 Gem; standard unpolarized star HD 37124) yields a polarization of ∼ 2.6 ± 0.8 percent (polarization angle 60^{o} ± 1^{o}). Interstellar polarization was checked via field stars close to the host galaxy but produced low results (almost comparable with the instrumental polarization of 0.3 ± 0.8 percent).”

2000 November 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green